I need to know how to solve the following problems:
3/4n+16=2-1/8n
3/2y-y=4+1/2y
1/4(7-3g)=g/8
18-3.8x+7.36-1.9x
2[st3(s-1)]+18

Answer by checkley77(12844)

3/4n+16=2-1/8n
3n/4+n/8=2-16
(2*2n+n)/8=-14
5n/8=-14
5n=-14*8
5n=-112
n=-112/5
n=-22.4 answer.
-----------------------------------------
3/2y-y=4+1/2y
3y/2-y=4+y/2
3y/2-y-y/2=4
(3y-2y-y)/2=4
0y/2=4 no solution for y.
-----------------------------------------
I'll leave the rest for you to practice on.
1/4(7-3g)=g/8
18-3.8x+7.36-1.9x
2[st3(s-1)]+18